OpenStudy (anonymous):

solve p=2(l+w) for l.

OpenStudy (anonymous):

L=(p-w)/2

OpenStudy (anonymous):

P= 2(l+w) P=2l+2w P-2w=2l Divide both sides by two. P-w=l I'm sorry I made a mistake the first time...